# SlimLM: An Efficient Small Language Model for On-Device Document Assistance

### v2

@@ Abstract
While small language models (SLMs) show promises for mobile deployment, their real-world performance and applications on smartphones remains underexplored. We present SlimLM, a series of SLMs optimized for document assistance tasks on mobile devices. Through extensive experiments on a Samsung Galaxy S24, we identify the optimal trade-offs between model size (ranging from 125M to 7B parameters), context length, and inference time for efficient on-device processing. SlimLM is pre-trained on SlimPajama-627B and fine-tuned on DocAssist, our constructed dataset for summarization, question answering and suggestion tasks. Our smallest model demonstrates efficient performance on S24, while larger variants offer enhanced capabilities within mobile constraints. We evaluate SlimLM against existing SLMs, showing comparable or superior performance and offering a benchmark for future research in on-device language models. We also provide an Android application, offering practical insights into SLM deployment. Our findings provide valuable insights and illuminate the capabilities of running advanced language models on high-end smartphones, potentially reducing server costs and enhancing privacy through on-device processing.
@@ 1 Introduction
Our results show that SlimLM models perform comparably or even better than existing SLMs of similar sizes across standard metrics such as BLEU Papineni et al. (2002), ROUGE Lin (2004), Semantic Textual Similarity (STS), Self-BLEU Zhu et al. (2018) for text diversity and GEval Liu et al. (2023). The smallest model SlimLM-125M demonstrates efficient performance on S24, making it suitable for widespread deployment. Larger variants, up to 1B parameters, offer enhanced capabilities while still operating within mobile constraints. To demonstrate real-world applicability, we develop an Android application showcasing SlimLM’s document assistance capabilities (Sec. 4).
@@ 2.2.1 Data Collection
We utilize our proprietary tools to compile a diverse collection of 82,850 publicly available documents, primarily consisting of illustrations, presentation slides, and spreadsheets. This dataset also includes machine-generated documents to ensure a comprehensive representation of various document types. We extract the document contents and prepare them for pre-processing to ensure the data is suitable for model fine-tuning.
@@
We employ Tiktoken OpenAI (2023b) in conjunction with GPT-3.5-turbo to tokenize the documents. Each document is segmented into 5 chunks, with each chunk containing a maximum of 200 tokens. This segmentation ensures that the maximum number of tokens per document after pre-processing is 1,000. Consequently, documents with fewer than 1,000 tokens remain unaltered, while longer documents are truncated. Table 2 presents the statistical analysis of token distribution per document, including the mean, standard deviation, and range of token counts, both before and after pre-processing.
@@
Table 2: Statistical comparison of token distribution per document before and after pre-processing 82,850 documents. The table shows the mean $\pm$ standard deviation and the range of token counts for each processing stage.
@@ 2.2.2 Data Annotation
Table 4 provides insight into the token usage statistics for the commercial LLM in annotating 82,850 documents. The relatively low standard deviation in completion tokens suggests consistent-length responses across different documents, which is desirable for maintaining annotation quality and consistency. The annotation process yields $\sim$414K examples for DocAssist. Of these, $\sim$2K examples were randomly selected for the test set, with the remaining examples allocated to the training set.
@@
Table 4: Token usage statistics for the commercial LLM in annotating 82,850 documents.
@@ 4 Use Case
Figure 1: Loading MobileLLM paper Liu et al. (2024) and interacting with AI assistant without internet access.
@@
When a document is loaded, such as a legal contract, the app instantly generates a summary, suggests relevant questions, and provides quick answers to user queries, all without internet connectivity. This streamlined process allows professionals to grasp essential information rapidly and identify areas needing closer examination while maintaining document confidentiality and improving overall user experience. Users can also interact with the document by chatting with the AI assistant (Fig. 1).
@@ 6 Conclusion
In this work, we introduce SlimLM models optimized for document assistance tasks. We identify the optimal balance between model size, inference time, and maximum context length for efficient processing on real mobile devices. Our specialized DocAssist dataset, constructed from $\sim$83K documents, enabled fine-tuning of SlimLM for three critical document assistance tasks. SlimLM models, ranging from 125M to 1B parameters, demonstrate comparable or superior performance to existing SLMs of similar sizes across standard metrics, while efficiently handling up to 800 context tokens. To showcase real-world applicability, we develop an Android application featuring SlimLM’s document assistance capabilities, paving the way for widespread deployment of efficient, on-device language models for enhanced user privacy and reduced server costs.

### v3

@@ Abstract
While small language models (SLMs) show promises for mobile deployment, their real-world performance and applications on smartphones remains underexplored. We present SlimLM, a series of SLMs optimized for document assistance tasks on mobile devices. Through extensive experiments on a Samsung Galaxy S24, we identify the optimal trade-offs between model size (ranging from 125M to 7B parameters), context length, and inference time for efficient on-device processing. SlimLM is pre-trained on SlimPajama-627B and fine-tuned on DocAssist, our constructed dataset for summarization, question answering and suggestion tasks. Our smallest model demonstrates efficient performance on S24, while larger variants offer enhanced capabilities within mobile constraints. We evaluate SlimLM against existing SLMs, showing comparable or superior performance and offering a benchmark for future research in on-device language models. We also provide a research demo, offering practical insights into SLM deployment. Our findings provide valuable insights and illuminate the capabilities of running advanced language models on high-end smartphones, potentially reducing server costs and enhancing privacy through on-device processing.
@@ 1 Introduction
Our results show that SlimLM models perform comparably or even better than existing SLMs of similar sizes across standard metrics such as BLEU Papineni et al. (2002), ROUGE Lin (2004), Semantic Textual Similarity (STS), Self-BLEU Zhu et al. (2018) for text diversity and GEval Liu et al. (2023). The smallest model SlimLM-125M demonstrates efficient performance on S24, making it suitable for widespread deployment. Larger variants, up to 1B parameters, offer enhanced capabilities while still operating within mobile constraints. To demonstrate real-world applicability, we develop a research demo showcasing SlimLM’s document assistance capabilities (Sec. 4).
@@ 2.2.1 Data Collection
We utilize our proprietary tools to compile a diverse collection of documents, primarily consisting of illustrations, presentation slides, and spreadsheets. This dataset also includes machine-generated documents to ensure a comprehensive representation of various document types. We extract the document contents and prepare them for pre-processing to ensure the data is suitable for model fine-tuning.
@@
We employ Tiktoken OpenAI (2023b) to tokenize the documents. Each document is segmented into 5 chunks, with each chunk containing a maximum of 200 tokens. This segmentation ensures that the maximum number of tokens per document after pre-processing is 1,000. Consequently, documents with fewer than 1,000 tokens remain unaltered, while longer documents are truncated. Table 2 presents the statistical analysis of token distribution per document, including the mean, standard deviation, and range of token counts, both before and after pre-processing.
@@
Table 2: Statistical comparison of token distribution per document before and after pre-processing the documents. The table shows the mean $\pm$ standard deviation and the range of token counts for each processing stage.
@@ 2.2.2 Data Annotation
Table 4 provides insight into the token usage statistics for the commercial LLM in annotating the documents. The relatively low standard deviation in completion tokens suggests consistent-length responses across different documents, which is desirable for maintaining annotation quality and consistency. The annotation process yields $\sim$414K examples for DocAssist. Of these, $\sim$2K examples were randomly selected for the test set, with the remaining examples allocated to the training set.
@@
Table 4: Token usage statistics for the commercial LLM in annotating the documents.
@@ 4 Use Case
When a document is loaded, such as a legal contract, the app instantly generates a summary, suggests relevant questions, and provides quick answers to user queries, all without internet connectivity. This streamlined process allows professionals to grasp essential information rapidly and identify areas needing closer examination while maintaining document confidentiality and improving overall user experience. Users can also interact with the document by chatting with the AI assistant.
@@ 6 Conclusion
In this work, we introduce SlimLM models optimized for document assistance tasks. We identify the optimal balance between model size, inference time, and maximum context length for efficient processing on real mobile devices. Our specialized DocAssist dataset, constructed from $\sim$83K documents, enabled fine-tuning of SlimLM for three critical document assistance tasks. SlimLM models, ranging from 125M to 1B parameters, demonstrate comparable or superior performance to existing SLMs of similar sizes across standard metrics, while efficiently handling up to 800 context tokens. To showcase real-world applicability, we develop a research demo featuring SlimLM’s document assistance capabilities, paving the way for widespread deployment of efficient, on-device language models for enhanced user privacy and reduced server costs.